OpenStudy (anonymous):

A social club charters a bus at a cost of $2805 to take a group of members on an excursion to Atlantic City. At the last minute, 4 people in the group decide not to go. This raises the transportation cost per person by $4. How many people originally intended to take the trip?

OpenStudy (anonymous):

2805 = x * y 2805 = (x-4) * (y+4)

OpenStudy (anonymous):

^He's right

OpenStudy (anonymous):

x=y x^2 = 2805 i would say

OpenStudy (anonymous):

but im not sure

OpenStudy (anonymous):

^How would that be right?

OpenStudy (anonymous):

The first set of equations is correct. Solve for x and y. x is the number of people and y is the cost per person

OpenStudy (anonymous):

not sure how to do so

OpenStudy (anonymous):

How many people originally intended to take the trip?!

OpenStudy (anonymous):

tejas?

OpenStudy (anonymous):

x=55 and y=51. Originally, 55 people intended to go on the trip.